Calculate the definite integral y = 𝑑 ^3 + 2𝑒 ^ 4𝑑 + 1/ 𝑑 + 3 sin 2𝑑)


1
Expert's answer
2022-05-25T15:57:23-0400
"y=\int (𝑡 ^3 + 2𝑒 ^ {4t} + \dfrac{1}{t} + 3 \sin (2𝑡))dt"

"=\dfrac{t^4}{4}+\dfrac{1}{2}e^{4t}+\ln|t|-\dfrac{3}{2}\cos(2t)+C"